A lawn(rectangular shape is to have an area of 86M.Forty metres of timber edging is available to complaetely edge the lawn.determine the lawn length and width so that all timber edging is used
Answer by checkley71(8403)

X*Y=86 & 2X+2Y=40 THUS 2X=40-2Y OR X=20-Y THEN SUBSTITUTE (20-Y) FOR X IN THE FIRST EQUATION WE GET (20-Y)Y=86 OR 20Y-Y^2-86=0 OR -Y^2+20Y-86=0 OR
Y^2-20Y+86=0 USING THE QUADRATIC EQUATION (-B+/-SQRT(B^2-4AC))/2A WE GET ANSWERS OF:
-10+/-SQRT14 OR 13.7442 & 6.258